OpenStudy (anonymous):

derivative of ∫(-2,sin(x)) (cos(t^5)+t) dt

OpenStudy (amistre64):

\[\int_{a}^{b}f(t)~dt=F(b)-F(a)\] \[\frac{d}{dx}\left[\int_{a}^{b}f(t)~dt\right]=\frac{d}{dx}\left[F(b)-F(a)\right]\] \[\frac{d}{dx}\left[\int_{a}^{b}f(t)~dt\right]=\frac{d}{dx}[F(b)]-\frac d{dx}F(a)]\] \[\frac{d}{dx}\left[\int_{a}^{b}f(t)~dt\right]=f(b)b'-f(a)a'\]

OpenStudy (amistre64):

since this is a general setup, it proves that we can go from the top to the bottom; so we really dont have to worry about the middle stuff ....

OpenStudy (amistre64):

since the derivative of -2 is 0 ... all it amounts to really is:\[f(b)b'\]

OpenStudy (anonymous):

That makes sense, thanks!